If y varies inversely with x, solve y if the constant of variation k=6 and k= 1/3
Please help!!!
Thanks in advance!

Answer by Boreal(15235)

y=k/x
if k=6, y=6/x
if k=1/3, y=k/(1/3)x=3k/x
Usually with these problems, you are given y and x and find k. Then you are asked to find y given a different x. The k found in the first part is used to solve the second.